OpenStudy (anonymous):

Prove that \[\tan(x+y) = \frac{\tan(x) + \tan(y)}{1 - \tan(x) \tan(y)}\]

OpenStudy (anonymous):

use the addition formulas for sin and cos.

OpenStudy (anonymous):

\[\frac{\sin(X+y)}{\cos(X+Y)}\] =\[\frac{sixcosy+cosxsiny}{cosxcosy-sinxsiny}\] dividing numerator and denominator by cosxcosy we get \[\frac{tanx+tany}{1-tanxtany}\]